# Can LLMs Generate Tabular Summaries of Science Papers? Rethinking the Evaluation Protocol

### v2

@@ 2 Related Works
Prior works primarily attempt to generate scientific tables through two stages: schema induction and value extraction. For schema induction, early methods like entity-based table generation Zhang and Balog (2018) focused on structured input, while recent work has explored schema induction from user queries Wang et al. (2024) and comparative aspect extraction Hashimoto et al. (2017). For value extraction, various approaches such as document-grounded question-answering Kwiatkowski et al. (2019); Dasigi et al. (2021); Lee et al. (2023), aspect-based summarization Ahuja et al. (2022), and document summarization DeYoung et al. (2021); Lu et al. (2020) have been proposed to extract relevant information. Beyond these methods, several datasets have been introduced to support scientific table-related tasks, such as TableBank Li et al. (2020), SciGen Moosavi et al. (2021), and SciTabQA Lu et al. (2023). Recently, Newman et al. (2024) proposed streamlining schema and value generation with LLMs sequentially and curated a large-scale benchmark for evaluation. However, all these methods assume a clean and fully relevant set of papers and rely on predefined captions or abstract-based schemas, which risk missing key details. In contrast, we argue for an evaluation approach where candidate papers include tangentially relevant or distracting papers, aligning more closely with real-world literature review workflows.
@@ 4.1 Constructing User Demand Prompts
To verify that our collected user demands align with our objective, we visualize: (1) the distribution of the number of tokens in the original and modified user demands, and (2) the ratio of captions and user demands of different lengths that have token overlap with the schema or values. From Figure 2, we observe that our modified user demands are generally longer than the original captions, providing a more detailed description of the table’s goal. Furthermore, as shown in Table 1, user demands exhibit a significantly lower overlap ratio with the schema and table values, resulting in fewer overlapping tokens. This ensures a fairer subsequent evaluation.
@@ 4.2 Paper Retrieval Simulation
After selecting these candidates, we conduct human annotations to verify whether they should indeed be excluded from the table. Given that annotating these tables requires expert knowledge in computer science, we recruit seven postgraduate students with research experience in the field as annotators. To ensure they are well-prepared for the task, the annotators undergo rigorous training, including pilot annotation exams. Their task is to make a binary decision on whether a given distractor paper—based on its title, abstract, user demand, the ground-truth table, and the titles and abstracts of all referenced papers—should be included in the table. Each table contains annotations for 10 papers, with each distractor paper initially assigned to two randomly selected annotators. If both annotators agree on the label, it is finalized. Otherwise, two additional annotators review the paper until a consensus is reached. In the first round, the inter-annotator agreement (IAA) is 94% based on pairwise agreement, and the Fleiss’ Kappa Fleiss (1971) score is 0.73, indicating a substantial level of agreement Landis and Koch (1977). Finally, for each table, we randomly select a number of distractor papers between $[m,10]$ and merge them with $R$ to form $C$.
@@ 4.3 Evaluation via LLM-based Utilization
We introduce three key aspects for evaluating a table in terms of its usability: (1) Schema: whether a specific column is included in the generated schema, (2) Unary Value: whether a particular cell from the ground-truth table appears in the generated table, and (3) Pairwise Value: whether relationships between two cells remain consistent in the generated table.
@@
We guide GPT-4o in generating these binary QA pairs based on the ground-truth table. For the first two aspects, we generate QA pairs for all columns and cells, whereas for the third aspect, we randomly sample 10 pairs of cells per table and synthesize them into QA pairs. We then prompt GPT-4o to answer these questions based on the generated table, providing yes/no responses. If the answer cannot be found, the model is instructed to respond with “no,” and vice versa for “yes.” The ratio of “yes” answers indicates how well the generated table preserves the schema, individual values, and pairwise relationships. This represents the recall of the ground-truth table, measuring how much original information is retained in the generated table.
@@ 6.1 Experiment Setup
To demonstrate the generalizability of our method and evaluations, we conduct experiments using two proprietary and three open-source LLMs as backbone model representatives: GPT-4o OpenAI (2024b), GPT-4o-mini OpenAI (2024a), DeepSeek-V3 (685B; DeepSeek-AI et al., 2024), LLAMA-3.3 (70B; Dubey et al., 2024), and Mistral-Large (123B; Mistral-AI, 2024). We apply all baseline methods and our proposed method to each model and use our evaluation framework to assess the quality of the generated tables based on our benchmark, focusing on four aspects: paper selection (Paper), schema content overlap (Schema), single-cell value overlap (Unary Value), and comparisons across cells (Pairwise Value). For paper selection, we use recall as the metric to measure the number of ground-truth papers successfully selected [^1] Recall is emphasized instead of precision since missing relevant papers is more critical than including slightly noisy ones in literature review tasks.. For the latter three tasks, we report precision (P), recall (R), and F1 scores (F1), as explained in §4.3.
@@ 6.2 Main Evaluation Results
(3) While unary values are well preserved, pairwise comparisons suffer substantial losses. Most methods, especially our proposed approach, extract unary values with relatively high F1 scores. However, extracting and maintaining pairwise relationships remains challenging. For instance, using LLAMA-3.3, our method achieves a unary F1 score of 51.1 but drops to 41.0 for pairwise values. This trend is consistent across different models, suggesting that while individual entries are correctly identified, capturing the relationships between them remains difficult. The significant gap highlights the challenge of preserving complex relational comparisons within the generated tables.
@@ 6.4 Validation of Utilization-Based Evaluation
Table 4: Comparison between GPT-4o and human annotators on 300 QA pairs per method. We report the proportion of “yes” answers by each and their overall agreement.
6.5 Computational Cost Comparisons
Method Success Rate #Tokens Avg. Runtime
Baseline 1 48.2% 128K 37
Baseline 2 98.2% 167K 118
Newman et al. 99.7% 110K 208
Ours 100.0% 118K 194

### v3

@@ 2 Related Works
Prior works primarily attempt to generate scientific tables through two stages: schema induction and value extraction. For schema induction, early methods like entity-based table generation Zhang and Balog (2018) focused on structured input, while recent work has explored schema induction from user queries Wang et al. (2024) and comparative aspect extraction Hashimoto et al. (2017). For value extraction, various approaches such as document-grounded question-answering Kwiatkowski et al. (2019); Dasigi et al. (2021); Lee et al. (2023), aspect-based summarization Ahuja et al. (2022), and document summarization DeYoung et al. (2021); Lu et al. (2020) have been proposed to extract relevant information. Beyond these methods, several datasets have been introduced to support scientific table-related tasks, such as TableBank Li et al. (2020), SciGen Moosavi et al. (2021), and SciTabQA Lu et al. (2023). Recently, Newman et al. (2024) proposed streamlining schema and value generation with LLMs sequentially and curated a large-scale benchmark for evaluation. However, all these methods assume a clean and fully relevant set of papers and rely on predefined captions or abstract-based schemas, which risk missing key details. In contrast, we argue for an evaluation approach where candidate papers include tangentially relevant or distracting papers, aligning more closely with real-world literature review workflows Padmakumar et al. (2025).
@@ 4.1 Constructing User Demand Prompts
To verify that our collected user demands align with our objective, we visualize: (1) the distribution of the number of tokens in the original and modified user demands, and (2) the ratio of captions and user demands of different lengths that have token overlap with the schema or values. From Figure 2, we observe that our modified user demands are generally longer than the original captions, providing a more detailed description of the table’s goal. Furthermore, as shown in Table 1, user demands exhibit a significantly lower overlap ratio with the schema and table values, resulting in fewer overlapping tokens.
@@ 4.2 Paper Retrieval Simulation
After selecting these candidates, we conduct human annotations to verify whether they should indeed be excluded from the table. Given that annotating these tables requires expert knowledge in computer science, we recruit seven postgraduate students with research experience in the field as annotators. To ensure they are well-prepared for the task, the annotators undergo rigorous training, including pilot annotation exams. Their task is to make a binary decision on whether a given distractor paper—based on its title, abstract, user demand, the ground-truth table, and the titles and abstracts of all referenced papers—should be included in the table. Each table contains annotations for 10 papers, with each distractor paper initially assigned to two randomly selected annotators. If both annotators agree on the label, it is finalized. Otherwise, two additional annotators review the paper until a consensus is reached. In the first round, the inter-annotator agreement (IAA) is 94% based on pairwise agreement, and the Fleiss’ Kappa Fleiss (1971) score is 0.73, indicating a substantial level of agreement Landis and Koch (1977). Finally, for each table, we randomly select a number of distractor papers between $[m,10]$.
@@ 4.3 Evaluation via LLM-based Utilization
We introduce three key aspects for evaluating a table in terms of its usability: (1) Schema: whether a specific column is included in the generated schema, (2) Unary Value: whether a particular cell from the ground-truth table appears in the generated table, (3) Pairwise Value: whether relationships between two cells remain consistent in the generated table.
@@
We guide GPT-4o in generating binary QA pairs based on the ground-truth table. For the first two aspects, we generate QA pairs for all columns and cells, whereas for the third, we randomly sample 10 cell pairs per table and synthesize them into QA pairs. We then prompt GPT-4o to answer these questions based on the generated table, providing yes/no responses. If the answer cannot be found, the model is instructed to respond with no,” and vice versa for yes.” The ratio of “yes” answers indicates how well the generated table preserves the schema, individual values, and pairwise relationships. This represents the recall of the ground-truth table, measuring how much original information is retained in the generated table.
@@ 6.1 Experiment Setup
To demonstrate the generalizability of our method and evaluations, we conduct experiments using two proprietary and three open-source LLMs as backbone model representatives: GPT-4o OpenAI (2024b), GPT-4o-mini OpenAI (2024a), DeepSeek-V3 (685B; DeepSeek-AI et al., 2024), LLAMA-3.3 (70B; Dubey et al., 2024), and Mistral-Large (123B; Mistral-AI, 2024). We apply all baseline methods and our proposed method to each model and use our evaluation framework to assess the quality of the generated tables based on our benchmark, focusing on four aspects: paper selection (Paper), schema content overlap (Schema), single-cell value overlap (Unary Value), and comparisons across cells (Pairwise Value). For paper selection, we use recall as the metric to measure the number of ground-truth papers successfully selected. For the latter three tasks, we report precision (P), recall (R), and F1 scores (F1), as explained in §4.3.
@@ 6.2 Main Evaluation Results
(3) While unary values are well preserved, pairwise comparisons suffer substantial losses. Most methods, especially our proposed approach, extract unary values with relatively high F1 scores. However, extracting and maintaining pairwise relationships remains challenging. This trend is consistent across different models, suggesting that while individual entries are correctly identified, capturing the relationships between them remains difficult. The significant gap highlights the challenge of preserving complex relational comparisons within the generated tables.
@@ 6.4 Validation of Utilization-Based Evaluation
Table 4: Comparison between GPT-4o and human annotators on 300 QA pairs. We report the proportion of “yes” answers by each and their overall agreement.
@@ 7 Conclusion
Acknowledgments
We thank the JHU CLSP community for discussions and inspiration and HKUST KnowComp community for helping with data annotations. The authors of this paper were supported by the NSFC Fund (U20B2053) from the NSFC of China, the RIF (R6020-19 and R6021-20), the GRF (16211520 and 16205322) from RGC of Hong Kong and ONR grant (N0001424-1-2089).
@@ A.2 Evaluation Implementations
A.3 Computational Cost Comparisons
Method Success Rate #Tokens Avg. Runtime
Baseline 1 48.2% 128K 37
Baseline 2 98.2% 167K 118
Newman et al. 99.7% 110K 208
Ours 100.0% 118K 194